OpenStudy (anonymous):

8x^3+4x^2-32x=112 can smeone plz show me how to solve this

OpenStudy (rajee_sam):

8x³ + 4x² - 32x -112 = 0 4( 2x³ + x² - 8x - 28) = 0 2x³ + x² - 8x -28 = 0 Now we have to factor this

OpenStudy (rajee_sam):

Do you know how to do it?

OpenStudy (anonymous):

the original problem was (5/x+2)-(3/x+2)=12/x^2-4, I followed the original steps you showed me, but I think I did something wrong

OpenStudy (anonymous):

I am stuck at ((3(x+2)+5(x+2)/(x+2)(x+2))=3x+6+5x+10/x^2+4=8x+16/x^2+4=12/x^2-4 now what's next?

OpenStudy (rajee_sam):

so the original problem had the same denominator for both fractions on the LHS. ( x+2) So you just subtract. (5-3) / (x + 2) = 12 / (x² -4) 2 / (x+2) = 12 / (x² -4) now on the RHS the denominator you have to factor it. it is of the form (a² - b² ). So the factors will be (a + b ) ( a-b) so here our factors will be ( x +2 ) and (x-2) 2/(x+2) = 12 / [(x+2)(x-2)] Now you multiply both sides by (x+2) since you have that factor in the denominator on both sides. 2 = 12 / (x-2) now cross multiply 2 (x-2) = 12 2x -4 = 12 2x = 16 x = 8

OpenStudy (rajee_sam):

with these type problems you may not encounter same steps.

OpenStudy (rajee_sam):

you have to remember basic rules

OpenStudy (rajee_sam):

1. Factor quadratic functions if any and cancel out common factors. 2. combine fractions after taking LCD. 3. Cross multiply fractions on either sides. 4.Simplify terms by combining like terms.

OpenStudy (rajee_sam):

@aprilsages

OpenStudy (anonymous):

got it sorry I was just reading all of it

OpenStudy (anonymous):

ok got it thanks so much!

OpenStudy (rajee_sam):

the first mistake you did was doing (x+2) (x+2) for LCD. both denominators are same. So you don't have to do that

OpenStudy (anonymous):

yeah I found that out once u told me, I was sorta skeptical about it in the first place, it felt wrong

OpenStudy (rajee_sam):

It is not required for this problem but still I noticed something in your work which is not right (x +2 ) (x +2 ) is not x² + 4 

OpenStudy (anonymous):

ooooooo

OpenStudy (anonymous):

k got it!